OpenStudy (anonymous):

A mouse has made holes in opposite corners of a rectangular kitchen. Starting from its hole in the northwest corner, the mouse scurries 4 meters along the length of the kitchen to reach a piece of cheese in the southwest corner. Then the mouse scurries 6 meters along the width of the kitchen to its other hole in the southeast corner. Finally the mouse scurries back to the first hole. What is the total distance the mouse scurries? If necessary, round to the nearest tenth.

OpenStudy (retireed):

If you draw the picture, I bet you can solve the problem.

OpenStudy (anonymous):

15 meters.

OpenStudy (anonymous):

If the shortest side is 4 meters and then the length of the rectangular kitchen is 6 meters from corner to corner, then the length of the side would have to be smaller than 6 but bigger than four and since there is nothing to say that there are any decimals, it's length is five meters and the total distance 15.

OpenStudy (dumbcow):

 using pythagorean thm \[x^2 = 4^2 + 6^2\] \[x = \sqrt{16+36} = \sqrt{52} = 2 \sqrt{13}\] distance mouse traveled was: \[10 + 2\sqrt{13}\]